# An image speaks a thousand words, but can everyone listen? On image transcreation for cultural relevance

### v2

@@ 5 Related Work
Image-editing models have evolved over the years from being capable of single editing tasks like style transfer Gatys et al. (2015, 2016) to handling multiple such tasks in one model Isola et al. (2017); Choi et al. (2018); Huang et al. (2018); Ojha et al. (2021). Today, their capabilities range from performing targeted editing that preserves spatial layout, local in-painting, and most notably, following natural language instructions Brooks et al. (2023). We choose, InstructPix2Pix Brooks et al. (2023) to experiment with since their model allows users to give natural language instructions, as opposed to other models requiring text labels, captions, segmentation masks, example output images and so on. The model is capable of following a wide variety of instructions, ranging from concrete ones like swap sunflowers with roses to abstract ones like make it Paris or make it the 1900s. [^9] https://www.timothybrooks.com/instruct-pix2pix From a practical standpoint, this model is consistently one of the most downloaded image-editing models on HuggingFace, [^10] https://huggingface.co/models?pipeline_tag=image-to-image&sort=downloads indicating its wide usage in the community.
@@ 6 Conclusion
In this paper, we introduce the task of image transcreation with machine learning systems, where we culturally adapt visual content to suit a target audience. Translation has traditionally been limited to language, but with increased consumption of multimedia content, translating all modes in a coherent way is essential. We build three pipelines comprising state-of-the-art generation models, and show that end-to-end image editing models are incapable of understanding cultural contexts, but using LLMs and retrievers in the loop helps boost performance. We create a challenging two-part evaluation dataset: (i) concept which is simple, cross-culturally coherent, and diverse; and (ii) application which is curated from education and stories. We conduct an extensive human evaluation and show that even the best models can only translate 5% images for select countries (like Nigeria) in the easier concept dataset and no image translation is successful for some countries (like Portugal) in the harder application dataset. Our code and data will be released to facilitate future work in this new, exciting line of research.
@@ Appendix A Example Outputs
Here, we include sample outputs from the pipelines for select images. All three pipelines have their own set of limitations, indicating that we have a long way to go before we can solve this task. Limitations and patterns observed for each pipeline can be found below:
@@
The models seem to associate flags and colors in them with a particular country/culture and includes these features in the edited images irrespective of the objects mentioned in the caption prompts. Some examples can be seen at Figure 12, where the American flag colors are applied over the Burger to make it relevant to the United States. Similarly, Figure 16 includes Brazil map and flag as part of the editing process.
@@
Both the image editing models preserve spatial dimensions, as can be seen in Figure 21
@@
The obtained images through the retrieval pipeline seem to be noisy with a low precision but high recall. Some of the images are better representatives of that country’s culture compared to the other two pipelines, given that they are real images.
@@
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
Figure 15: Application: Education; Target: Japan — Task: count the number of cherries. Here, even though there is a semantic drift from cherries to flowers, the output of cap-edit can be still be used to solve the task.
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
@@ Appendix D Continued analysis of human evaluation
Q3: spatial-layout – For e2e-instruct and cap-retrieve, we observe similar trends as those observed in Q1). For cap-edit, while it scores mid to high on visual changes, it surprisingly maintains spatial layout, performing similar to e2e-instruct. This signifies that even though cap-edit makes visual edits, it does so while preserving spatial layout, helpful for audiovisual translation like in Doraemon, Inside Out and so on.
Q5: naturalnesscap-retrieve receives highest scores here since these are natural images retrieved from the internet. cap-edit receives a significant number of 4s, because it doesn’t look as natural as retrieved images, but probably natural enough, as discussed in §A.2.
Q6: offensiveness – Almost no images are found to be offensive, which is encouraging.

### v3

@@ 5 Related Work
Cultural diversity in image generation: Several recent works investigate cultural awareness of text-to-image (T2I) systems typically highlighting biases towards certain cultures. Hutchinson et al. (2022) highlight how under-specified prompts show gender and western cultural biases, Jha et al. (2024) analyse regional stereotypical markers in generated images, Naik and Nushi (2023) discuss occupational biases of neutral prompts and personality trait associations with limited groups of people, Cho et al. (2023) reveal skin-tone biases and Bird et al. (2023) discuss associated risks of these biases for society. Some other works focus on ways to probe for and evaluate cultural relevance of generated images. Ventura et al. (2023) derive prompt templates to unlock the cultural knowledge in T2I systems, and Hall et al. (2023) evaluate the realism and diversity of T2I systems when prompted to generate objects from across the world. While all of these works are targeted towards assessing and mitigating cultural biases in pre-trained models, our work is targeted towards an application (i.e. transcreating visual content) that would benefit by such efforts that improve the cultural understanding and diversity of image generation models.
Image-editing models have evolved over the years from being capable of single editing tasks like style transfer Gatys et al. (2015, 2016) to handling multiple such tasks in one model Isola et al. (2017); Choi et al. (2018); Huang et al. (2018); Ojha et al. (2021). Today, their capabilities range from performing targeted editing that preserves spatial layout, local in-painting, to edits that can follow natural language instructions Brooks et al. (2023). We choose InstructPix2Pix Brooks et al. (2023) to experiment with, given its flexibility to prompt with natural language instructions, as opposed to other models requiring text labels, captions, segmentation masks, example output images and so on. It has also consistently been one of the most downloaded image-editing models on HuggingFace. [^10] https://huggingface.co/models?pipeline_tag=image-to-image&sort=downloads As discussed in Section 4 however, these models are only capable of making color, shape and style changes, and lack a deeper understanding of natural language. No image-editing works have tackled the semantically complex task of cultural transcreation. We hope that our work paves the way to building image-editing models that truly understand natural language, which can benefit multiple applications, including ours.
@@ 6 Conclusion
In this paper, we introduce a new task of image transcreation with machine learning systems, where we culturally adapt visual content to suit a target audience. Translation has traditionally been limited to language, but with increased consumption of multimedia content, translating all modes in a coherent way is essential. We build three pipelines comprising state-of-the-art generation models, and show that end-to-end image editing models are incapable of understanding cultural contexts, but using LLMs and retrievers in the loop helps boost performance. We create a challenging two-part evaluation dataset: (i) concept which is simple, cross-culturally coherent, and diverse; and (ii) application which is curated from education and stories. We conduct an extensive human evaluation and show that even the best models can only translate 5% images for select countries (like Nigeria) in the easier concept dataset and no image transcreation is successful for some countries (like Portugal) in the harder application dataset. Our code and data is released to facilitate future work in this new, exciting line of research.
@@ Appendix A Example Outputs
Here, we include sample outputs from the pipelines for select images. All pipelines have their own set of limitations, indicating that we have a long way to go before we can solve this task. Patterns observed for each pipeline can be found below:
@@
The models seem to associate flags and colors in them with a particular country/culture and includes these features in the edited images irrespective of the objects mentioned in the caption prompts. Some examples can be seen in Figure 16, where the American flag colors are applied over the Burger to make it relevant to the United States. Similarly, Figure 19 includes Brazil map and flag as part of the editing process. The code to run this pipeline is here. [^12] https://anonymous.4open.science/r/image-translation-6980/src/pipelines/e2e-instruct.py We simply pass in the original image with the instruction make this image culturally relevant to COUNTRY.
@@
Spatial dimensions are highly preserved in this pipeline as can be seen in Figure 25. This can sometimes lead to undesirable outcomes or the outputs to look unnatural, as shown below. The code to run this pipeline can be found here. [^13] https://anonymous.4open.science/r/image-translation-6980/src/pipelines/caption-llm_edit.py. The exact prompts used for captioning and LLM editing can be found in Table D.2.
(a) original
(b) cap-edit
Figure 8: Example of how preserving the spatial layout of the original image can lead to unnatural looking outputs. Here, the final image shows a cup of chai, but a typical cup of chai looks different in India.
@@
The obtained images through the retrieval pipeline seem to be noisy with a low precision but high recall. Some of the images are better representatives of that country’s culture compared to the other two pipelines, given that they are real images. However, this pipeline also suffers from failure cases of retrieving images which may be too different from the source image or retrieving irrelevant outputs. Examples are shown below:
(a) original
(b) cap-retrieve
Figure 9: Example of how the retrieved output may at times look completely different from the original image.
(a) original
(b) cap-retrieve
Figure 10: Example of how the retrieved output may be irrelevant/noisy. Here, we can see it behaving like a bag-of-words since the llm-edit used to prompt for retrieval is: A sunflower stands tall against the backdrop of a clear blue sky in India.
A.4 GPT4-o + GPT-4 + DALLE-3
We use the GPT-4 family of models for this pipeline. Since DALLE-3 works with detailed prompts Betker et al. (2023), we prompt GPT4-o to give detailed captions for images. We use GPT4 to edit these captions and prompt DALLE-3 to generate images. To make the images look natural, we add "photo, photograph, raw photo, analog photo, 4k, fujifilm photograph" to the prompt. [^14] https://www.reddit.com/r/dalle/comments/1au10g6/generate_realistic_pictures_with_dalle/ Even then, the images do have a distinct style. Qualitatively, we observe that the captions and caption-edits capture fine-grained details which shorter captions in the previous two pipelines cannot. The overall pipeline can be found in Figure 12. All visualizations can be found in the released code repository. Note that GPT4-o + DALLE-3 outputs could not be human evaluated since their APIs were released on May 13, 2024. Further, the images’ distinct style defeats the purpose of randomizing pipeline outputs for human evaluation.
@@ A.4 GPT4-o + GPT-4 + DALLE-3
Figure 12: Pipeline for GPT4-based experiments.
@@ B.4.7 USA
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
@@ Appendix D Continued analysis of human evaluation
C0: visual-change – First we ask whether the image has been edited at all, to help understand if the edits make sense in the questions that follow. Across all countries, cap-retrieve maximally edits images, with roughly 90% scoring 5 (Figure 6). This is expected since here the original image is not input at all in producing the final image. e2e-instruct on the other hand makes no edit sometimes, with 40-60% images being given a score of 1. For countries like Brazil and US, this pipeline overwhelmingly paints the image with the flag or flag colors (§A.1), explaining the relatively lower number of 1s.
C1: semantic-equivalence – Here, we ask that if an edit is made (C0$<3$) is it a meaningful one? In Figure 6, we observe that cap-edit scores highest, while cap-retrieve’s performance varies based on the country (lower for countries with low digital presence).
C2: spatial-layout – For e2e-instruct and cap-retrieve, we observe similar trends as those observed in Q1). For cap-edit, while it scores mid to high on visual changes, it surprisingly maintains spatial layout, performing similar to e2e-instruct. This signifies that even though cap-edit makes visual edits, it does so while preserving spatial layout, helpful for audiovisual translation like in Doraemon, Inside Out and so on.
C4: naturalnesscap-retrieve receives highest scores here since these are natural images retrieved from the internet. cap-edit receives a significant number of 4s, because it doesn’t look as natural as retrieved images, but probably natural enough, as discussed in §A.2.
C5: offensiveness – Almost no images are found to be offensive, which is encouraging.
Figure 23: Human ratings for the application dataset: Our goal is to test whether the edited image can be used for the application as before (E/S1), and whether it increases cultural relevance (E/S2). We plot the count of images that can do both above (E/S1+E/S2), and observe that even the best pipeline cannot transcreate any image successfully in some cases, like for Portugal.
D.1 Application Dataset
D.2 Quantitative Metrics
For image-editing, these typically capture how closely the edited image matches – (i) the original image; and (ii) the edit instruction. Following suit, we calculate two metrics: a) image-similarity: we embed the original image and each of the generated images using DiNO-ViT Caron et al. (2021) and measure their cosine similarity; and b) country-relevance: we embed the text – This image is culturally relevant to {COUNTRY}, and the edited images using CLIP Radford et al. (2021) and calculate their cosine similarity. We present results for both metrics in Figures 20 and 21. A discussion on correlation of these metrics with human evaluation is in §C.
We find that overall for image-similarity, e2e-instruct scores highest, closely followed by cap-edit, while cap-retrieve lags behind, consistent with human ratings. For the country-relevance score, we observe a similar trend as that for C3: cultural-relevance.
@@ D.2 Quantitative Metrics
Prompts used for the pipelines described in Section 2. InstructBLIP Prompt (Captioning)
Concept Dataset
"A short image description:"
Application Dataset (Education)
"This image is from a math worksheet titled: TASK. Describe the image such that it talks about details relevant to the task of the worksheet. The output should be ONLY ONE sentence long."
Application Dataset (Stories)
"This image is from a storybook for children. Caption the image such that it describes details relevant to the story."
GPT3.5 Prompt (LLM-editing)
Concept Dataset
"Edit the input text, such that it is culturally relevant to COUNTRY. Keep the output text of a similar length as the input text. If it is already culturally relevant to COUNTRY, no need to make any edits. The output text must be in English only.
Input: "
Application Dataset (Education)
"Edit the input text, such that it is culturally relevant to COUNTRY. The text describes an image in a math worksheet titled: TASK. Hence, make sure the edit preserves the intent of the task in the worksheet. Keep the output text to be of a similar length as the input text. If it is already culturally relevant to COUNTRY, there is no need to make any edits. The output text must be in English only.
Input: "
Application Dataset (Stories)
"Edit the input text, such that it is culturally relevant to COUNTRY. The text describes an image in a storybook for children. Make sure the edit preserves the meaning of the story. Keep the output text to be of a similar length as the input text. If it is already culturally relevant to COUNTRY, there is no need to make any edits. The output text must be in English only.
Input: "